Fermi National Laboratory


2008 Accelerator Update Archive

December 22, 2008 - December 29, 2008
- Eight stores provided ~108 hours of luminosity
- Experts changed power amplifiers for Linac RF stations # 3 & 5
- Store 6683 set New Record with a luminosity of 360.95E30
- Pelletron VME problems
- Stores 6684 and 6688 lost due to quenches
- Stuck Kautzky valve reseated twice
- TeV sector A3 harmonic filter and VCB work

The Integrated Luminosity for 12/22/08 to 12/29/08 was 59.42 Pb-1.

December 19, 2008 - December 22, 2008
- Four stores provided ~48 hours of luminosity
- Pbar set New stacking record with a rate of 28.32 mA/hr
- Recycler cooling problems fixed
- Service building problems due to low temperatures
- MiniBooNE, NuMI, and Recycler suffer from kicker problems

The Integrated Luminosity for 12/15/08 to 12/22/08 was 47.59 Pb-1.

December 15, 2008 - December 17, 2008
- Three stores provided ~31.75 hours of luminosity
- Booster kicker work continues
- T988 ends their run
- Store 6647 lost - B1 quench

December 12, 2008 - December 15, 2008
- Four stores provided ~58.5 hours of luminosity
- NuMI resumes taking beam
- Booster kicker problems resolved
- CAMAC crate restart resolves ECOOL problem
- The Integrated Luminosity for 12/8/08 to 12/15/08 was 73.73 Pb-1. This set a new record for weekly integrated luminosity

December 10, 2008 - December 12, 2008
- three stores provide ~39 hours of luminosity.
- Pbar sets new stacking record.
- Booster quad problem.

December 8, 2008 - December 10, 2008
- Four stores provided ~41.25 hours of luminosity
- Linac resolves I- Source sparking and quadrupole problems
- Experts repair seized F27 water skid pump

December 5, 2008 - December 8, 2008
- Last week's integrated luminosity was 55.29 Pb-1
- Four stores provided ~46.5 hours of luminosity
- Store 6619 aborted, there was no quench
- Store 6625 set New Record Luminosity: 355.07E30
- TeV sector A3 wet engine repaired

December 3, 2008 - December 5, 2008
- Three stores provided ~20.75 hours of luminosity
- Operations aborted store 6609 just before a TeV quench
- Pbar kicker problems fixed

December 1, 2008 - December 3, 2008
- Three stores provided ~34 hours of luminosity
- Cryo system Cold box problem fixed
- Store 6605 - second best luminosity: 339.19E30
- Store 6606 - third best luminosity: 335.6E30

November 24, 2008 - December 1, 2008
- Nine stores provided ~143 hours of luminosity
- Store 6593 sets New Record* with a luminosity of 344.8E30
- I- Source brought online
- Network storm requires many system reboots
- Booster vacuum problems fixed
- Booster GMPS problems
- CHL cold box lost

*Integrated Luminosity for 11/24/08 to 12/1/08 = 63.65 Pb-1 (This set a New Record; previous best was 57.54 Pb-1)

November 21, 2008 - November 24, 2008
- Five stores provided ~59.25 hours of luminosity
- Booster backup LCW return pump brought online
- MTA takes commissioning beam
- MI LCW 3-way valve problems
- Store 6589 set 2nd highest luminosity

November 19, 2008 - November 21, 2008
- Three stores provided ~22.55 hours of luminosity
- A LCW leak in the Linac gallery was found and repaired
- TeV quench lost store 6578
- Stuck Kautzky valve repaired

November 17, 2008 - November 19, 2008
- Three stores provided ~35.25 hours of luminosity
- Store 6566 lost due to quench
- The #2 NuMI horn might need replacing

November 14, 2008 - November 17, 2008
- Four stores provided 53.55 hours of luminosity
- Booster Chopper repaired
- NuMI horn off
- LRF5 problems

November 12, 2008 - November 14, 2008
- Two stores provided ~34.7 hours of luminosity
- Operations and Controls expert resolve database errors
- Linac experts discover phase errors causing trips
- TeV quench near end of store setup
- Cryo experts fix stuck Kautzky valve

November 10, 2008 - November 12, 2008
- TeV turn on completed
- Two stores provided ~13 hours of luminosity
- Pelletron brought back online

November 7, 2008 - November 10, 2008
- TeV continues to recover
- D2 wet engine gets overhaul
- Pelletron power supply replaced
- Next store might come Tuesday midnight shift

November 5, 2008 - November 7, 2008
- One store provided ~6.5 hours of luminosity
- Fermi suffers from unexpected 345 kV power outage
- Recovery in progress
- Next store? Maybe on Monday, 11/10/08

November 3, 2008 - November 5, 2008
- Troubled D0 cryo system temporarily hooked to TeV
- Store 6541 sets New Record with a luminosity of 332.12E30
- Linac has RF station problems

October 31, 2008 - November 3, 2008
- Antiproton target replaced
- Store 6535 set New Record a with luminosity of 318.98E30
- Linac experts resolve KRF3 troubles

October 29, 2008 - October 31, 2008
- Four stores provided ~42.5 hours of luminosity
- Electron Cooling restored
- Store 6529 aborted
- CUB Column 1 brought online
- Accumulator has stacking problem

October 27, 2008 - October 29, 2008
- Two stores provided ~23 hours of luminosity
- Linac experts repair KRF7
- TeV suffers two quenches
- Pelletron off due to GVN problems

October 24, 2008 - October 27, 2008
- four stores provide ~ 57.25 hours of luminosity
- Main Injector power supply problems
- Kicker causes a quench in the Tevatron
- Klystron problems in the Linac

October 22, 2008 - October 24, 2008
- Four stores provided ~26.5 hours of luminosity
- Cryo experts repair stuck Kautzky valve
- LRF3 power amplifier replaced
- Store 6507 aborted: CAMAC power supply failed

October 20, 2008 - October 22, 2008
- Three stores provided ~42.25 hours of luminosity
- Klystron Vernier flow turbine replaced
- MI-50 Upper power supply brought back online
- Pbar dry engine flywheel replaced

October 17, 2008 - October 20, 2008
- four stores provided ~ 61 hours and 15 minutes of luminosity
- Tevatron abort
- Main Injector power supply problems
- Linac gradient problems

October 15, 2008 - October 17, 2008
- Three stores provided ~31.5 hours of luminosity
- New MTest experiment T977 (MINERvA) begins
- Recycler stash (~200E10) lost during card change
- TeV conducts beam-beam compensation study

October 13, 2008 - October 15, 2008
- Four stores provided ~36 hours of luminosity
- Accumulator and NuMI kickers fixed
- Pelletron resumes providing cooling

October 10, 2008 - October 13, 2008
- Two stores provided 12.5 hours of luminosity
- Shutdown ends and startup begins
- Pelletron problems continue
- Lithium Lens trip

October 8, 2008 - October 10, 2008
- The shutdown ends and startup begins
- Linac will begin sending beam Friday afternoon
- The TeV cryo systems should be cold by late Friday night

October 6, 2008 - October 8, 2008
- The accelerator complex is shutdown for a week of maintenance and repair
- The shutdown work is on schedule
- Startup is scheduled to begin on Friday

October 3, 2008 - October 6, 2008
- Fours stores provided ~21.5 hours of luminosity
- TeV conducts crystal collimator studies
- TeV quench caused by cold compressor trip
- LCW leaks in Booster, MI, and TeV enclosures found
- One week shutdown begins

October 1, 2008 - October 3, 2008
- Three stores provided ~38.75 hours of luminosity
- Electron Cooling GVM fails
- Booster LCW pumps put in manual control
- Recycler takes bigger, but fewer stashes

September 29, 2008 - October 1, 2008
- Five stores provided ~26.5 hours of luminosity
- MI LCW column at CUB regenerated
- TeV quench during scrape
- Unplanned sector B4 power outage

September 26, 2008 - September 29, 2008
- Five stores provided ~48.5 hours of luminosity
- MI-30 power supply transformer problems
- TeV experts investigate low beta quenches
- Pelletron suffers from node problems

September 24, 2008 - September 26, 2008
- Three stores provided ~32.5 hours of luminosity
- TeV quench during shot setup
- TeV sector A3 flywheel changed
- TeV thermal scans reveals hot switch

September 22, 2008 - September 24, 2008
- Two stores provided ~27 hours and 19 minutes of luminosity
- TeV quenched during shot setup
- MIRF watchdog trips

September 19, 2008 - September 22, 2008
- Four stores provided ~54 hours and 26 minutes of luminosity
- TeV sector F0 vacuum burst cascades to A1 transfer line
- TeV quenched due to bad IO crate power supply

September 17, 2008 - September 19, 2008
- Three stores provided ~39 hours and 19 minutes of luminosity
- MTA survey work
- LRF5's PA replaced
- Velocity meter problems resolved

September 15, 2008 - September 17, 2008
- Two stores provided ~32 hours and 10 minutes of luminosity
- LCW leak discovered on Pbar magnet
- Power supply fault causes loss of stack
- Inspection of TeV power supply knife-switches begins
- TeV quench diagnosed: Collimator at fault

September 12, 2008 - September 15, 2008
- Three stores provided ~49 hours and 55 minutes of luminosity
- Pbar's lithium lens replaced
- Rain, rain, rain: heroic efforts from many people kept lab operating
- Linac water skid #3 conductivity probe fails

September 10, 2008 - September 12, 2008
- Three stores provided 43 hours and 24 minutes of luminosity
- Kautz Road substation repaired
- Linac Klystron RF station 7 repaired

September 8, 2008 - September 10, 2008
- Four stores provided ~34 hours and 22 minutes of luminosity
- Computer room power glitch causes quench
- Raccoon shorts out Kautz Road substation

September 5, 2008 - September 8, 2008
- Fours stores provided 66 hours and 26 minutes of luminosity
- Lithium Lens repaired
- MIRF has modulator problems

September 3, 2008 - September 5, 2008
- four stores provided ~ 35 hours and 30 minutes of luminosity
- Problems with the Pbar lithium lens
- Booster RF problems

September 1, 2008 - September 3, 2008
- Three stores provided ~37 hours and 23 minutes of luminosity
- H- chopper problem fixed
- Two TeV quenches - store 6393 lost

August 29, 2008 - September 1, 2008
- Fours stores provided 65 hours and 45 minutes of luminosity
- Pbar's pulse magnet replaced
- LRF3's LCW leak repaired
- H- Source's output adjusted
- MIRF5 causes problems for NuMI

August 27, 2008 - August 29, 2008
- Three stores provided ~21 hours and 23 minutes
- Recycler installs amplifier on for 2-4 stochastic cooling
- TeV shutdown on Thursday for CDF maintenance work
- TeV quench during wet squeeze - B0Q6 power supply
- Pbar Pulsed magnet bad, needed replacement.

August 25, 2008 - August 27, 2008
- Four stores provided ~40 hours and 38 minutes of luminosity
- Controls still searching for the cause of QF12 problems
- MTest T984 completes their experimental run
- CDF offline due to solenoid problems
- Run Coordinator schedules 10-hour access for CDF on Thursday

August 22, 2008 - August 25, 2008
- Four stores provided 54 hours and 1 minute of luminosity
- LRF4 repaired
- F4 wet engine overhauled
- B3 bypass switch replaced
- Controls experts working on QF12 problems

August 20, 2008 - August 22, 2008
- Two stores provided 34 hours and 54 minutes of luminosity
- Lithium Lens gradient voltage raised to 2600 volts
- C4 cold compressor replaced
- Timing jitter found on QF107 and QF12

August 18, 2008 - August 20, 2008
- Three stores delivered ~40 hours and 2 minutes
- Booster experts resolved their kicker (MP02) problems
- The TeV quenched with no beam
- Pbar experts make adjustments to stochastic cooling
- MTest experiment T979 completes its run

August 15, 2008 - August 18, 2008
- Four stores provided ~61 hours and 10 minutes of luminosity
- Bad MTest magnet replaced
- Pbar has beam valve and beam permit problems
- Booster MP02 kicker magnet problems
- SciBooNE completes their data taking

August 13, 2008 - August 15, 2008
- three stores provided ~ 43 hours and 45 minutes of luminosity
- Transformer failure on the Pbar lens
- Ground fault on MT5E magnet string

August 8, 2008 - August 13, 2008
- Six stores provided ~101 hours of luminosity
- Controls and kicker problems in Pbar
- ECool front end problems
- Network problems cause Operators to take the master switch
- Linac I- source problems

August 6, 2008 - August 8, 2008
- Three stores provided ~32 hours and 37 minutes of luminosity
- Booster kicker transformer replaced
- MI Center Anode power supply repaired
- TeV aborted - no cause found
- A0 CAMAC crate replaced - bad power supply
- Pbar Debuncher BPM system upgraded

August 4, 2008 - August 6, 2008
- Twos stores provided ~16 hours and 44 minutes of luminosity
- TeV recovered from Lightning strike on C3 service building
- TeV sector F2 power supply still a problem
- Power glitch trips off equipment, but stack, stash, and store remains intact

August 1, 2008 - August 4, 2008
- Two stores provided ~17 hours and 21 minutes of luminosity
- TeV D0Q6 power supply repaired
- Store quenched during ramp
- TeV sector D1 cold compressor replaced
- Store 6334 lost due to lightning strike to C3 service building

July 30, 2008 - August 1, 2008
- Three stores provided ~43 hours 15 minutes of luminosity
- Pbar magnet H717 repaired in place
- FESS installs pond water spray heads at MI-20

July 28, 2008 - July 30, 2008
- Three stores provided ~36 hours and 17 minutes of luminosity
- TeV sector D1 wet engine bearings replaced
- TeV tilt meters registers earthquakes
- Pbar magnet D:H717 develops leak and arcing
- Experts attempt to repair Pbar magnet in place

July 25, 2008 - July 28, 2008
- Four stores provided ~64 hours and 30 minutes of luminosity
- D0 experiences network problems due to a small power outage
- E-Cool frontend problems

July 23, 2008 - July 25, 2008
- Two stores provided ~25 hours and 45 minutes of luminosity
- Recycler stash lost due to failed relay
- MI finds cause of ground faults

July 21, 2008 - July 23, 2008
- One store provided ~27 hours and 2 minutes
- TeV experts hunt for illusive ground fault
- Stack lost due to broken A:LQ 480-volt breaker

July 18, 2008 - July 21, 2008
- Four stores provided ~47 hours and 57 minutes of luminosity
- Booster WAPS resistor replaced
- MI PSLOOP trips due to ground faults
- Accumulator power supply trip; stack lost
- 7/21/08 access to find TeV ground fault and for Pbar repairs

July 16, 2008 - July 18, 2008
- Two stores provided 28 hours and 52 minutes of luminosity
- Store 6300 terminated to prevent possible quench
- NuMI repairs closed loop cooling problem
- Shutdown for MI heat exchanger work

July 14, 2008 - July 16, 2008
- Three stores provided ~31 hours and 40 minutes of luminosity
- NuMI finds air leak on horn closed loop
- 480-volt breaker replaced for Bartleson low beta power supply
- TeV sector A4 wet engine fly wheel replaced
- Shutdown at 6 AM, Thursday 7/16/08

July 11, 2008 - July 14, 2008
- Three store provided ~15 hours and 50 minutes of luminosity
- Two stores lost due to quench and abort
- MI MECAR problems
- MTest SEM failed
- NuMI resumes taking beam
- Experts search for TeV ground fault

July 09, 2008 - July 11, 2008
- Two store provided ~22 hours and 30 minutes of luminosity
- Losses in Tevatron delay shot set-up
- NuMI close to taking data once again

July 07, 2008 - July 09, 2008
- One store provided ~24 hours and 40 minutes of luminosity
- Kautz Rd substation trips due to harmonic filter failure
- Pbar raises LCW temperature regulation
- Pbar beam stop motor replaced
- FESS repairs Booster East Gallery cooling
- Recycler stash lost - bulk power supply trip

July 02, 2008 - July 07, 2008
- Eight stores provided ~80 hours and 18 minutes of luminosity
- TeV sector C4 LCW temperature high alarm due to valve problem
- TeV heat exchanger work completed
- NuMI takes beam for target scans
- LRF3 modulator LCW leak repaired
- Store 6266 sets New Record Luminosity with 318.49E30

June 30, 2008 - July 02, 2008
- Fours stores provided ~42 hours and 6 minutes of luminosity
- TeV abort problem fixed by TEVCAL reboot
- Booster Watchdog trips fixed - switched to backup WavTek generator

June 27, 2008 - June 30, 2008
- Four stores provided ~ 63 hours and 3 minutes of luminosity
- Water group prepares TeV for next round of heat exchanger cleaning
- Experts put TeV sector C3 cryo system into JT mode to adjust wet engine

June 25, 2008 - June 27, 2008
- Three stores provided ~37 hours of luminosity
- Problems with Booster Watchdog trips resolved
- TeV/MI experts conduct injection studies

June 23, 2008 - June 25, 2008
- Three stores provided ~34 hours and 25 minutes of luminosity
- Store 6241 lost due to quench
- FESS replaces B0 pump
- FESS drops Pbar 95 LCW temperature by 5 per request

June 20, 2008 - June 23, 2008
- Four stores provided ~62 hours and 38 minutes of luminosity
- TeV heat exchanger at B1 required cleaning
- MTest experiment T971 has reclaimed MT6
- Pbar suffers vacuum leak at AP-2

June 18, 2008 - June 20, 2008
- Three stores provided ~29 hours and 48 minutes of luminosity
- TeV quench during end-of-store study
- Experts test the MI LCW system

June 16, 2008 - June 18, 2008
- Three stores provided ~32 hours and 42 minutes of luminosity
- Booster GMPS taken offline for repairs
- The MTest T979 experiment completes its run
- Stash lost due to MI interlock trip

June 13, 2008 - June 16, 2008
- Four stores provided ~46 hours and 34 minutes of luminosity
- High temperatures cause problems
- Nitrogen Dewar at A0 develops leak
- Sunday morning lightning strikes cause store 6221 to quench
- Lightning strikes trip off many devices

June 11, 2008 - June 13, 2008
- Three stores provided ~43 hours and 47 minutes of luminosity
- New old style target installed in Pbar
- D0 interaction point changed
- New Meson MTest experiment (T979) begins
- Rising temperatures begin causing problems

June 9, 2008 - June 11, 2008
- Three stores provided ~35 hours and 23 minutes of luminosity
- FESS cleans MI pond water strainers
- Pbar Debuncher kicker problems
- Pbar Target stops rotating: 8-10 hours access begins Thursday morning

June 6, 2008 - June 9, 2008
- Five stores provided ~47 hours and 58 minutes of luminosity
- The New MTest experiment (T971) began taking beam
- Store 6203: fourth highest luminosity store - 288.8E30
- Sore 6207 aborted due to lightning strike

June 4, 2008 - June 6, 2008
- Three stores provided 21 hours and 43 minutes of luminosity
- Linac experts repair KRF6
- TeV vacuum and Cryo problems fixed
- D0 problem requires 8-hour access
- Store 6200 initial luminosity 298.41E30

June 2, 2008 - June 4, 2008
- Four stores provided 33 hours and 17 minutes of luminosity
- NuMI Horn cooling system fails; Horn replacement begins
- Pbar sets New Stacking Record: 27.01 mA/hr
- I- Source water resistor flushed

May 30, 2008 - June 2, 2008
- Five stores provided 61 hours and 33 minutes of luminosity
- NuMI horn had LCW flow problem
- Main Injector timing problem
- New Stacking Record: 26.91mA per hour

May 28, 2008 - May 30, 2008
- Three stores provided 37 hours and 16 minutes of luminosity
- All Booster RF stations are up and running
- Cross Gallery Computer Room suffers from chilled water leak

May 26, 2008 - May 28, 2008
- Two stores provided 29 hours and 50 minutes of luminosity
- F23 Yard Breaker problems
- NuMI horn LCW problem temporarily resolved
- TeV quench during shot setup
- MTest T978 experiment done - back in July

May 23, 2008 - May 26, 2008
- Four stores provided 44 hours and 1 minute of luminosity
- Pbar sets new stacking record with a rate of ~26.6 per hour
- TeV quench during shot setup
- NuMI horn #1 LCW problem halts experiment

May 21, 2008 - May 23, 2008
- Two stores provided 38 hours and 43 minutes of luminosity
- The Run Coordinator reported that nearly all the necessary jobs were completed during this shutdown
- NuMI LCW leak found and fixed
- Antiproton target replaced in record time
- Store 6154 aborted

May 19, 2008 - May 21, 2008
- Three stores provided 39 hours and 1 minute of luminosity
- Store 6150 aborted due to a power supply problem
- Thursday, 5/22/08, Everything except the TeV and Pbar Rings will be shutdown for maintenance

May 16, 2008 - May 19, 2008
- Five stores provided 65 hours and 51 minutes of luminosity
- Short TeV shutdown on Friday for necessary repairs
- TeV quench at F4 with no beam
- Linac experts resolve quadrupole problems
- Pbar LCW leak found and repaired
- Last week's integrated luminosity was 55.03 inverse picobarns

May 12, 2008 - May 14, 2008
- Three stores provided 32 hours and 58 minutes of luminosity
- Linac Klystron RF station 7 repaired

May 9, 2008 - May 12, 2008
- Five stores provided 62 hours and 34 minutes of luminosity
- Linac experts investigate Klystron trips
- Controls experts replace MI-31 network switch
- MI and Booster experts work on NuMI intensity
- 30mA stack lost
- Record integrated luminosity of 56.085 inverse picobarns last week

May 7, 2008 - May 9, 2008
- Three stores provided 29 hours and 5 minutes of luminosity
- Several quenches and aborts troubled the TeV
- Booster repairs two RF stations during access

May 5, 2008 - May 7, 2008
- Two stores provided 44 hours and 6 minutes of luminosity
- Quick action resolves C0 pond water pump failure
- Pbar implements a faster mode for stashing
- TeV experts conduct end-of-store studies

May 2, 2008 - May 5, 2008
- five stores provided 49 hours and 30 minutes of luminosity
- Two power glitches
- Bad cooling pump at C0

April 30, 2008 - May 2, 2008
- Two stores provided 14 hours and 39 minutes of luminosity
- 4/30/08 Accelerator accesses for maintenance
- TeV antiproton kicker power supply trip
- 50E10 of Recycler stash lost
- Potential store lost

April 25, 2008 - April 30, 2008
- Two stores provided 35 hours and 40 minutes of luminosity
- The new Meson MTest T978 experiment begins taking beam
- Booster and MI RF fixed
- CDF and D0 access 4/30/08 for detector repairs

April 25, 2008 - April 28, 2008
- Three stores provided 46 hours and 32 minutes of luminosity
- Store quenched just before being established
- TeV experts worked on orbits
- Vacuum experts replace two TeV turbo stations

April 23, 2008 - April 25, 2008
- Two stores provided 14 hours and 7 minutes of luminosity
- Recycler stash lost
- Two TeV quenches
- MI heat exchangers successfully flushed
- Stashing efficiency low

April 21, 2008 - April 23, 2008
- Three stores provided 34 hours and 42 minute of luminosity
- A fourth store was established, but quenched immediately
- Pbar switches to rapid transfers

April 11, 2008 - April 21, 2008
- Six stores provided luminosity during this last week
- TeV experts found CDF aperture restriction that caused several quenches
- Booster ground faults caused by cables
- P2 vacuum burst delays D0 and CDF data taking
- Fermi monitors West Salem, IL earthquakes
- Power glitch trips off devices in Linac, Booster, and Accumulator

April 9, 2008 - April 11, 2008
- There were no stores established
- Preparation to replace Booster kicker MP02
- TeV suffered a multi house quench during shot setup
- Pelletron down for repairs
- The Accelerator Updates will stop for a week due to furlough

April 7, 2008 - April 9, 2008
- Two stores provided 34 hours and 31 minutes of luminosity
- Energy conservation mode used several times
- Booster kicker magnet (MP02) fails
- 10E10 of antiprotons lost
- NuMI and BNB beam halted to prepare fir Booster access
- Lightning strikes causes store to abort and TeV to quench
- On 4/8/08, Pbar set a new stacking record: 25.9mA/hr

April 4, 2008 - April 7, 2008
- Three stores provided 49 hours and 46 minutes of luminosity
- TeV A0 kicker repaired
- Recycler vector signal analyzer problems
- TeV turbo pump swapped
- Pelletron chiller problems
- Booster chopper cable repaired

April 2, 2008 - April 4, 2008
- One store provided 36 hours and 21 minutes of luminosity
- Support groups replace NuMI magnet
- Pbar installs rebuilt lithium lens
- D0 has front end problems
- Booster MP02 septa ground fault trips

March 31, 2008 - April 2, 2008
- Two stores provided 22 hours and 9 minutes of luminosity
- ComEd power glitch trips off Booster and Pbar equipment
- NuMI quadrupole magnet failed
- Pbar beam valves closings investigated
- Wednesday morning access to replace NuMI quadrupole magnet

March 28, 2008 - March 31, 2008
- Three stores provided 32 hours and 2 minutes of luminosity
- Pbar suffered from beam valve closures
- MI has vacuum failure in RF section
- Recycler stash aborted to access MI
- NuMI quadrupole magnet has internal LCW leak

March 26, 2008 - March 28, 2008
- Two stores provided 44 hours and 57 minutes of luminosity
- MI safety system testing completed
- Pbar lithium lens turbine replaced
- MINOS far detector power outage completed

March 24, 2008 - March 26, 2008
- Three stores provided 43 hours and 14 minutes of luminosity
- Modified Linac Spectrometer power supply gets remote controls
- Scheduled power outage for Sudan Mine on 3/27/08

March 21, 2008 - March 24, 2008
- Two stores provided 36 hours of luminosity
- Safety system testing done except for MI
- TeV quenched between stores
- Cryo experts repair B2 wet engine
- Linac and Booster experts repair Spectrometer

March 19, 2008 - March 21, 2008
- Two stores provided 38 hours and 30 minutes of luminosity
- MI cooling problems

March 17, 2008 - March 19, 2008
- Two stores provided 45 hours and 59 minutes of luminosity
- MI LCW put in recirculating mode
- MI-8 BPM problems fixed

March 14, 2008 - March 17, 2008
- Four stores provided 54 hours and 59 minutes of luminosity
- CUB column #1 need regeneration
- Store 5986 had a luminosity of 285.05E30, it was lost due to a quench after 21 hours
- Store 5989 set a NEW RECORD luminosity of 315.55E30
- Last weeks integrated luminosity a NEW RECORD with ~42.471 inverse picobarns

March 12, 2008 - March 14, 2008
- Four stores provided 27 hours and 10 minutes of luminosity
- Store 5977 quenched due to compressor trip
- TeV auto orbit smooth program working
- MI collimator studies improve stashing and TeV transfers

March 10, 2008 - March 12, 2008
- Three stores provided 33 hours and 28 minutes of luminosity
- Preaccelerator switched to I- Source
- Experts investigating AP1 transfer line trips
- NuMI resumed taking beam

March 7, 2008 - March 10, 2008
- Four stores provided 42 hours and 18 minutes of luminosity
- NuMI resumes using beam

March 5, 2008 - March 7, 2008
- One store provided 17 hours and 11 minutes of luminosity
- Pbar resumes stacking
- Safety system tests beginning
- NuMI will resume taking beam on Saturday, 3/8/08

March 3, 2008 - March 5, 2008
- No stores were established during this time period
- Pbar experts continue work on lithium lens
- MI replaces Q603 magnet
- TeV conditions antiproton kickers

February 29, 2008 - March 3, 2008
- Three stores provided 64 hours and 28 minutes of luminosity
- Experts work on lithium lens replacement
- Booster vacuum leak repaired
- Recycler kicker coolant vapor leak repaired
- Linac experts study momentum absorber vacuum problems

February 27, 2008 - February 29, 2008
- Two stores provided 16 hours and 30 minutes of luminosity
- Two Tev quenches
- Linac RF problems

February 25, 2008 - February 27, 2008
- Two stores provided 36 hours and 17 minutes of luminosity
- TeV experts conduct many studies
- Store 5934 hits third highest with luminosity of 286.4E30

February 15, 2008 - February 25, 2008
- Seven stores provided luminosity to the experiments during this period
- Experts repaired leak in Recycler kicker
- MTest experiment T953 completed their run
- Pbar sets record stack with 310mA of antiprotons

February 13, 2008 - February 15, 2008
- Three stores provided 31 hours and 39 minutes of luminosity
- Cryo system experts conduct wet engine maintenance
- MI collimators in use for TeV shots
- MTest taking beam
- Due to furlough, Accelerator Update on hiatus for one week

February 11, 2008 - February 13, 2008
- Three stores provided 34 hours and 42 minutes of luminosity
- Antiproton target moved to new position
- Pbar conducts Debuncher gain ramp studies
- MI to use collimators for next TeV shot
- New MTest user (T953) starts taking beam

February 8, 2008 - February 11, 2008
- Four stores provided 48 hours and 5 minutes of luminosity
- One store quenched before being established
- Booster found some melted wiring insulation
- TeV F13 vacuum pump repaired
- Pbar reported that they broke a record by stacking 2863mA of antiprotons last week

February 6, 2008 - February 8, 2008
- Three stores provided 39 hours and 15 minutes of luminosity
- Severe winds whip 345kV power lines around
- TeV suffers from vacuum problems

February 4, 2008 - February 6, 2008
- Two stores provided 37 hours and 33 minutes of luminosity
- TeV expert closes F4 valve, after pump failure, to preserve cryo system
- Switchyard experts prepare MTest beamline for next user
- MI repairs vacuum leak
- NuMI resets humidifier heaters

February 1, 2008 - February 4, 2008
- Three stores provided 43 hours and 21 minutes of luminosity
- TeV experts investigate two unusual quenches

January 30, 2008 - February 1, 2008
- Two stores provided 44 hours 51 minutes of luminosity
- MI successfully replaces bad quadrupole magnet (Q631)
- Cause of LRF3 trips found
- Jan. 2nd best month for luminosity with a total of 164.9 pb-1

January 28, 2008 - January 30, 2008
- Two stores provided 25 hours and 42 minutes of luminosity
- Pbar breaks goal of producing 25.19mA of antiprotons in one hour and sets new record
- SY experts tune beam to MTest
- Store 5872 lost due to quench at D0

January 25, 2008 - January 28, 2008
- Three stores provided 42 hours and 11 minutes of luminosity
- TeV corrector (T:HA24) power supply repaired
- Stash lost to either a MI quadrupole ground fault or a safety system fault
- Pelletron LCW cooling problem fixed
- 10 - 12 hour MI access scheduled for 6 AM on 1/30/08

January 23, 2008 - January 25, 2008
- Two stores provided 35 hours and 40 minutes of luminosity
- Pelletron experts trying to improve Electron Cooling
- Experts repaired Linac 800 MHz clock power supply
- Store 5861: initial luminosity of 270.05E30

January 21, 2008 - January 23, 2008
- Three stores provided 42 hours and 2 minutes of luminosity
- Pelletron experts continue to work on cooling
- Store 5860 - Luminosity of 243.85E30

January 18, 2008 - January 21, 2008
- Tevatron shrugs off problems to establish three new stores over chilly weekend.

January 16, 2008 - January 18, 2008
- Two stores provided 41 hours and 50 minutes of luminosity
- Store 5852 made it to the 10th spot for highest luminosity with a luminosity of 275.4E30

January 14, 2008 - January 16, 2008
- Three stores provided 19 hours and 9 minutes of luminosity
- The Recycler stashed antiprotons 22 times during this period

January 11, 2008 - January 14, 2008
- Three stores provided 57 hours 9 minutes of luminosity
- CDF solenoid had cryo system problem

January 9, 2008 - January 11, 2008
- Three stores provided 38 hours and 48 minutes of luminosity
- MI vacuum leak found and temporarily repaired
- MI implements 11-batch slipstack running mode

January 7, 2008 - January 9, 2008
- Three stores provided 28 hours and 37 minutes of luminosity
- Operators abort store 5832 to prevent quench
- New stacking record set with rate of 24.67E10/hr

January 1, 2008 - January 7, 2008
- Five stores were established
- Store 5830 has highest initial luminosity since the fall shutdown with 232.88E30
- TeV recovers quickly after magnet replacement

 
[an error occurred while processing this directive]

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.


Return to Current Accelerator Update

 
last modified 12/28/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory